# DynaNoise: Dynamic Probabilistic Noise Injection for Defending Against Membership Inference Attacks

### v1

@@ 5.2. Information Leakage in Deep Learning
MIDPUT Overall Privacy–Utility Trade-off metric.
$\text{MIDPUT}_{C}$, $\text{MIDPUT}_{L}$, $\text{MIDPUT}_{S}$ Per-attack MIDPUT metrics (Confidence, Loss, Shadow).
@@ 6. Proposed Approach
Our proposed approach dynamically mitigates membership inference attacks by adapting the noise injection process to the sensitivity of each individual query. The step-by-step processes of the proposed approach is provided in Algorithm 1. The system is composed of three primary components, as described below:
@@
    which quantifies the uncertainty of the prediction. We then define the sensitivity score $R(q)$ for query $q$ as:
@@
    ensuring that $R(q)\in[0,1]$. A higher $R(q)$ indicates that the model is highly confident (i.e., lower entropy) in its prediction, thus representing a greater risk for membership inference attacks.
@@
    Using the computed sensitivity score, we dynamically adjust the noise injected into the model’s output. We define the noise variance as a function of the sensitivity score:
@@
    where $\sigma_{0}^{2}$ is the base noise variance and $\lambda$ is a scaling parameter that amplifies the noise for higher-risk queries. The noise $\eta$ is then sampled from a Gaussian distribution:
    $\eta\sim\mathcal{N}(0,\sigma(q)^{2}),$
    and added to the raw output $f(q)$ of the model:
    $\tilde{f}(q)=f(q)+\eta.$
    This formulation ensures that queries with a higher sensitivity score receive proportionally more noise, thus effectively obfuscating any distinguishing membership signals.
@@
    To mitigate potential distortions introduced by noise injection while maintaining model accuracy, we apply a probabilistic smoothing operation by re-normalizing the perturbed outputs using a softmax function with a temperature parameter $T>1$:
    $\hat{f}(q)=\text{softmax}\left(\frac{\tilde{f}(q)}{T}\right).$
    The temperature $T$ controls the sharpness of the resulting probability distribution, providing a trade-off between smoothing and maintaining the discriminative power of the original output.
By integrating sensitivity analysis, dynamic noise injection, and probabilistic smoothing, our approach adapts to the context of each inference request. Hence, this specific defense mechanism effectively obfuscates membership signals in high-risk queries while maintaining overall data utility and predictive accuracy.
Figure 2. Overview of the proposed DynaNoise approach integrated into the AlexNet deep learning model for illustration. The noise injection process adapts to the model’s uncertainty, providing enhanced obfuscation against membership inference attacks.
@@
3:  Step1 - Compute Raw Output:
4:   Compute logits: $\mathbf{z}\leftarrow f(q)\in\mathbb{R}^{k}$.
5:   Compute probability vector: $\mathbf{p}\leftarrow\text{softmax}(\mathbf{z})$.
6:  Step2 - Sensitivity Analysis:
7:   Compute Shannon entropy: $H(\mathbf{p})\leftarrow-\sum_{i=1}^{k}p_{i}\log(p_{i})$.
8:   Set sensitivity score: $R(q)\leftarrow 1-\frac{H(\mathbf{p})}{\log k}$.
9:  Step3 - Dynamic Noise Injection:
10:   Compute noise variance: $\sigma(q)^{2}\leftarrow\sigma_{0}^{2}\,(1+\lambda\,R(q))$.
11:   Sample noise: $\boldsymbol{\eta}\sim\mathcal{N}(\mathbf{0},\,\sigma(q)^{2}\,I)$.
12:   Perturb logits: $\mathbf{\tilde{f}(q)}\leftarrow\mathbf{f(q)}+\boldsymbol{\eta}$.
13:  Step4 - Probabilistic Smoothing:
14:   Compute final output: $\hat{f}(q)\leftarrow\text{softmax}\Bigl{(}\frac{\mathbf{\tilde{f}(q)}}{T}\Bigr{)}$.
@@ 7. Evaluation
We evaluated our approach on several benchmark datasets and model architectures that are explained in Section 7.3. Our analysis focuses on three primary aspects: (i) the effectiveness of membership inference attacks; (ii) the impact on model accuracy (i.e., utility); and (iii) the computational cost of the defense mechanism.
@@ 7.1. Metrics
  • Membership Inference Defense Privacy–Utility Trade-off (MIDPUT): Our proposed metric for evaluating the trade-off between privacy and utility. Let
    $\Delta_{\text{acc}}=\text{test\_acc}_{\text{no\_def}}-\text{test\_acc}_{\text{def}},$
    and for each attack type $A\in\{\text{conf},\text{loss},\text{shadow}\}$,
    $\Delta_{A}=\text{attack\_acc}_{\text{no\_def}}^{(A)}-\text{attack\_acc}_{\text{def}}^{(A)}.$
    Then, the overall MIDPUT is defined as
    $\text{MIDPUT}=\left(\frac{\Delta_{\text{conf}}+\Delta_{\text{loss}}+\Delta_{\text{shadow}}}{3}\right)-\Delta_{\text{acc}},$
    and the per-attack MIDPUT metrics are given by
    $\text{MIDPUT}_{A}=\Delta_{A}-\Delta_{\text{acc}},\quad A\in\{\text{conf},\text{loss},\text{shadow}\}.$
    The MIDPUT metric is bounded within the range $[-1,1]$, where values closer to 1 represent strong privacy gains with minimal utility loss, and values approaching -1 indicate poor trade-offs with greater accuracy degradation than privacy improvement.
@@ 7.2. Membership Inference Attacks
We implemented three distinct membership inference attacks:
@@
    Shadow-Model Attack: Trains a shadow model on an auxiliary dataset drawn from the same distribution as the target model’s training dataset. From the shadow model’s outputs, features such as maximum confidence, cross-entropy loss, and confidence margin are extracted to train a binary classifier $A(\mathbf{x})$, where $\mathbf{x}$ is the feature vector. The membership decision is then based on the classifier’s output. In our setup, we allocate 70% of the available data to training and evaluation the target model, while the remaining 30% is reserved for constructing the shadow model and training the attack classifier. We use the same architecture for the shadow model as the target model to closely mimic its behavior, and we employ a logistic regression classifier as the final attack model.
Each membership inference attack is systematically evaluated under the following three defense conditions:
@@
    SELENA Defense: The SELENA defense approach is employed, which leverages ensemble learning and self-distillation to mitigate membership leakage prior to the execution of the attack.
@@ 7.4. Experimental Setup and Results
We conduct experiments on CIFAR-10, ImageNet-10, and SST-2 datasets. For each dataset, 70% of the data is used to train and test the target model, while the remaining 30% is allocated for training shadow and attack models. All models are trained for 15 epochs using stochastic gradient descent (SGD) with a learning rate of 0.01 and a batch size of 64. Experiments are conducted on a machine equipped with NVIDIA RTX 5000 GPU to ensure efficient training and evaluation.
@@
As our experimental baseline defense, we implement SELENA (Tang et al., 2022) as described in Section 2.2, using $K=25$ sub-models and $L=10$ partitions per sample. Each target model is evaluated against three types of membership inference attacks: Confidence Threshold Attack, Loss Threshold Attack, and Shadow Model Attack. For each attack, we report the attack success rate (ASR) under three conditions: (i) without any defense (None), (ii) after applying the SELENA defense mechanism, and (iii) after applying our proposed adaptive noise injection method (DynaNoise).
@@
To quantify the privacy–utility trade-off, we utilize our proposed metric called MIDPUT, as described in section 7.1. This metric captures the extent to which a defense reduces attack success rate while preserving model accuracy. We report both the overall MIDPUT and per-attack values ($\text{MIDPUT}_{C}$, $\text{MIDPUT}_{L}$, and $\text{MIDPUT}_{S}$) for each defense mechanism.
Table 3 summarizes the membership inference metrics and the corresponding MIDPUT values for each dataset and defense mechanism. For CIFAR-10, the baseline (no defense) model achieves a test accuracy of 0.8211, with ASR values of 0.6956 (Confidence), 0.7639 (Loss), and 0.7841 (Shadow). SELENA reduces these to 0.7668, 0.5446, 0.6307, and 0.6755, respectively, resulting in an overall MIDPUT of 0.0766. In contrast, DynaNoise maintains a high test accuracy of 0.8156 while significantly lowering the attack success rates to 0.2785 (Confidence), 0.4221 (Loss), and 0.5334 (Shadow). Consequently, the overall MIDPUT for DynaNoise is 0.331, with per-attack MIDPUT values of 0.4116, 0.3363, and 0.2452, respectively. This shows that DynaNoise clearly outperforms SELENA in reducing MIA success rate while preserving model accuracy on CIFAR-10.
A similar trend is observed on ImageNet-10: while the baseline test accuracy is 0.9165, SELENA degrades the accuracy significantly to 0.7085 and produces very low (even negative) MIDPUT values (overall MIDPUT of -0.0342). Conversely, DynaNoise preserves the test accuracy at 0.9115, reduces the attack metrics to 0.2248 (Confidence), 0.308 (Loss), and 0.5725 (Shadow), and achieves an overall MIDPUT of 0.3358. These results highlight that DynaNoise not only provides stronger privacy protection than SELENA but also avoids the substantial utility drop seen with SELENA on ImageNet-10.
On SST-2, The baseline DistilBERT model reaches a test accuracy of 0.8865. SELENA slightly increases the accuracy to 0.8945 but yields only marginal improvements in the attack metrics, resulting in an overall MIDPUT of 0.0392. In contrast, DynaNoise achieves a test accuracy of 0.8911 while significantly reducing the confidence attack ASR to 0.2354 and achieving a slightly better reduction in the shadow attack ASR, leading to an overall MIDPUT of 0.2091. Despite the already high baseline accuracy, DynaNoise achieves a more favorable trade-off compared to SELENA, offering meaningful privacy gains in the SST-2 NLP setting.
@@
Table 3. Model accuracy and MIA attack success rates along with the proposed MIDPUT metrics on different datasets under various defense mechanisms.

### v2

@@ 5.2. Information Leakage in Deep Learning
$\text{MIDPUT}_{Overall}$ Overall Privacy–Utility Trade-off metric.
$\text{MIDPUT}_{\text{Conf}},\text{MIDPUT}_{\text{Loss}}$ Per-attack MIDPUT metrics for Confidence-based and Loss-based attacks.
$\text{MIDPUT}_{\text{Shadow}},\text{MIDPUT}_{\text{LiRA}}$ Per-attack MIDPUT metrics for Shadow-model and LiRA attacks.
$\text{MIDPUT}_{\text{Ent}},\text{MIDPUT}_{\text{M-Ent}}$ Per-attack MIDPUT metrics for entropy-based and M-Entropy membership inference attacks.
@@ 6. Proposed Approach
Our proposed approach dynamically mitigates membership inference attacks by adapting the noise injection process to the sensitivity of each individual query. The complete procedure is summarized in Algorithm 1. The system is composed of three primary components, as described below.
@@
    which quantifies the uncertainty of the model’s prediction. We then define the normalized sensitivity score for query $q$ as:
@@
    ensuring $R(q)\in[0,1]$. A higher $R(q)$ indicates that the model is highly confident (i.e., low entropy) in its prediction, and therefore at greater risk of leaking membership information.
@@
    Using the computed sensitivity score, we dynamically adjust the noise added to the model’s output logits. The noise variance scales as:
@@
    where $\sigma_{0}^{2}$ is the base noise variance and $\lambda$ is a scaling parameter that amplifies the perturbation for higher-risk queries. Gaussian noise $\eta\sim\mathcal{N}(0,\sigma(q)^{2})$ is then added to the raw logits $f(q)$, producing perturbed outputs $\tilde{f}(q)=f(q)+\eta$. This ensures that high-risk queries receive proportionally more noise, effectively obfuscating membership signals.
@@
    To mitigate distortions introduced by noise injection while maintaining predictive performance, we apply a probabilistic smoothing operation by re-normalizing the perturbed logits using a temperature-scaled softmax:
    $\hat{f}(q)=\text{softmax}\!\left(\frac{\tilde{f}(q)}{T}\right),$
    where $T>1$ controls the sharpness of the probability distribution and provides a trade-off between smoothing and discriminative power.
By integrating sensitivity analysis, dynamic noise injection, and probabilistic smoothing, DynaNoise adapts to the context of each inference request. This design effectively obscures membership cues in high-risk queries while maintaining high accuracy and low computational overhead. Figure 2 provides an overview of the DynaNoise post-processing pipeline and its integration into the target model.
Figure 2. Overview of the proposed DynaNoise approach integrated into the AlexNet model for illustration. The noise injection process adapts to the model’s uncertainty, providing enhanced obfuscation against membership inference attacks.
@@
3:Step 1: Compute Raw Output
4:  Compute logits $f(q)\in\mathbb{R}^{k}$.
5:  Compute probability vector $p=\text{softmax}\!\bigl(f(q)\bigr)$.
6:Step 2: Sensitivity Analysis
7:  Compute Shannon entropy $H(p)=-\sum_{i=1}^{k}p_{i}\log(p_{i})$.
8:  Compute sensitivity score $R(q)=1-H(p)/\log k$.
9:Step 3: Dynamic Noise Injection
10:  Compute variance $\sigma(q)^{2}=\sigma_{0}^{2}\bigl(1+\lambda R(q)\bigr)$.
11:  Sample noise $\boldsymbol{\eta}\sim\mathcal{N}\!\bigl(\mathbf{0},\sigma(q)^{2}I_{k}\bigr)\in\mathbb{R}^{k}$.
12:  Perturb logits $\tilde{f}(q)=f(q)+\boldsymbol{\eta}$.
13:Step 4: Probabilistic Smoothing
14:  Compute final output $\hat{f}(q)=\text{softmax}\!\bigl(\tilde{f}(q)/T\bigr)$.
@@
Although DynaNoise is primarily an empirical defense, its adaptive noise modulation is theoretically motivated by the notion of sensitivity in Differential Privacy (DP). Here we explain how the query sensitivity measure $R(q)$ conceptually relates to DP and why this provides an interpretable foundation for privacy protection. Connection to differential privacy: In $(\epsilon,\delta)$-DP, the global sensitivity of a mechanism $f$ is defined as
$\Delta f=\max_{D,D^{\prime}}\|f(D)-f(D^{\prime})\|,$
where $D$ and $D^{\prime}$ differ by a single record. The Gaussian mechanism achieves privacy by adding noise proportional to this sensitivity, ensuring that the influence of any individual record on the output remains limited. In DynaNoise, the model’s output probability vector for a query $q$ serves as the observable function $f(D)$. Instead of computing dataset-level differences, DynaNoise estimates a per-query sensitivity proxy using Shannon entropy: queries with low entropy (high confidence) are empirically linked to greater membership risk. The normalized entropy complement $R(q)$ thus functions as a practical analogue to local sensitivity, determining the magnitude of noise required to conceal the query’s influence. Approximate privacy analogy: As defined in the previous section, DynaNoise scales the variance of the added Gaussian noise with $R(q)$. Since in DP the privacy loss parameter $\epsilon$ is inversely proportional to the noise magnitude, DynaNoise implicitly assigns each query an effective privacy budget:
$\epsilon(q)\propto\frac{1}{\sigma(q)}.$
Queries with higher sensitivity (low entropy) therefore receive larger perturbations, corresponding to smaller effective $\epsilon$ (stronger privacy), while low-sensitivity queries incur smaller noise to preserve utility.
6.1. Entropy as a Proxy for Sensitivity
DynaNoise modulates output noise according to the Shannon entropy $H(\mathbf{p}(q))$ of the model’s predictive distribution,
$\mathbf{p}(q)=\text{softmax}(\mathbf{z}(q)),\quad\mathbf{z}(q)=f_{\theta}(q).$
While entropy measures predictive uncertainty, it also serves as a lightweight surrogate for local sensitivity. Intuitively, entropy reflects how concentrated the model’s confidence is across classes, which determines how the output responds to small input perturbations. Formally, if $J_{f}(q)=\frac{\partial f_{\theta}(q)}{\partial q}$ denotes the output Jacobian, then under a first-order approximation $f_{\theta}(q+\delta)\approx f_{\theta}(q)+J_{f}(q)\delta$, and the resulting change in the softmax output scales with $\|J_{f}(q)\|$, a local Lipschitz proxy for sensitivity (Cisse et al., 2017; Novak et al., 2018).
As shown in prior works  (Ross and Doshi-Velez, 2018; Novak et al., 2018; Tsipras et al., 2019), low-entropy (peaked) predictions correspond to overconfident regions with sharper and more brittle decision boundaries, whereas high-entropy predictions yield smoother and more stable responses. Empirically, we observe a consistent relationship between entropy and Jacobian-based sensitivity across all evaluated models. Detailed scatter plots for CIFAR-10, ImageNet-10, and SST-2—showing Pearson correlations between $H(\mathbf{p}(q))$ and $\|J_{f}(q)\|_{2}$—are provided in Appendix A. These visualizations confirm that entropy reliably captures the model’s exposure level, supporting its use as a calibration-free proxy for sensitivity within DynaNoise.
@@ 7. Evaluation
We evaluated our approach on several benchmark datasets and model architectures that are explained in Section 7.3. Our analysis focuses on three primary aspects: (i) the effectiveness of membership inference defenses; (ii) the impact on model accuracy (i.e., utility); and (iii) the computational cost of the defense mechanism
@@ 7.1. Metrics
  • Membership Inference Defense Privacy–Utility Trade-off (MIDPUT): Our proposed metric for quantifying the trade-off between privacy protection and model utility. Let
    $\Delta_{\text{acc}}=\text{test\_acc}_{\text{no\_def}}-\text{test\_acc}_{\text{def}}$
    denote the accuracy drop induced by a defense, and for each attack $A\in\{\text{conf},\text{loss},\text{ent},\text{m-ent},\text{shadow},\text{lira}\}$, let
    $\Delta_{A}=\text{attack\_acc}_{\text{no\_def}}^{(A)}-\text{attack\_acc}_{\text{def}}^{(A)}$
    denote the corresponding reduction in attack success rate.
    The overall MIDPUT score is defined as
    $\text{MIDPUT}_{\text{Overall}}=\frac{1}{6}\sum_{A}\Delta_{A}-\Delta_{\text{acc}},$
    where the summation is taken over all evaluated attack types. The per-attack MIDPUT scores are given by
    $\text{MIDPUT}_{A}=\Delta_{A}-\Delta_{\text{acc}}.$
    Across our experiments, MIDPUT values typically lie in the range [1,1][-1,1], where values closer to $1$ indicate strong privacy gains with minimal utility loss, while values closer to 1-1 reflect unfavorable trade-offs dominated by accuracy degradation.
Extensibility of MIDPUT: The MIDPUT metric is not tied to any specific attack and can incorporate additional attacks in a modular way. In this work, we report both per-attack MIDPUT scores (e.g., $_{\text{Conf}}, MIDPUT$$_{\text{Loss}}$, $_{\text{Entropy}}, MIDPUT$$_{\text{M\text{-}Entropy}}$, $_{\text{Shadow}}, MIDPUT$$_{\text{LiRA}}$) as well as an overall MIDPUT that averages across attacks. This design serves two purposes: (i) it prevents conclusions from depending on a single attack model, and (ii) it enables future work to plug in stronger or newly proposed membership inference attacks without changing the structure of the evaluation protocol. In other words, MIDPUT is a general scoring framework for privacy–utility trade-offs rather than a metric tailored to any specific attack.
Interpretation and role of MIDPUT: Traditional privacy–utility evaluations often rely on Pareto frontiers, which visually illustrate trade-offs but are difficult to compare across models or attack settings. The proposed MIDPUT metric provides a complementary scalar summary that quantifies the joint balance between privacy preservation and utility retention. By aggregating the reduction in attack success rate and the preservation of accuracy across multiple attack types, MIDPUT yields a single interpretable score that enables consistent benchmarking and ranking of defenses.
[…]
@@ 7.2. Membership Inference Attacks
We implemented four distinct membership inference attacks:
@@
    Entropy-based attacks: Following Song and Mittal (Song and Mittal, 2021), we evaluate two entropy-based membership inference attacks that operate directly on the model’s output distribution. The Entropy attack infers membership by thresholding the Shannon entropy
    $H(\mathbf{p})=-\sum_{i}p_{i}\log p_{i},$
    with lower entropy indicating higher likelihood of membership. The Modified Entropy (M-Entropy) attack refines this criterion by incorporating the probability assigned to the true label, improving robustness across varying confidence profiles. Both attacks require only black-box access to predicted probabilities.
@@
    Shadow-model attack: Trains a shadow model on an auxiliary dataset drawn from the same distribution as the target model’s training dataset. From the shadow model’s outputs, features such as maximum confidence, cross-entropy loss, and confidence margin are extracted to train a binary classifier $A(\mathbf{x})$, where $\mathbf{x}$ is the feature vector. The membership decision is then based on the classifier’s output. In our setup, we allocate 70% of the available data to training and evaluating the target model, while the remaining 30% is reserved for constructing the shadow model and training the attack classifier. We use the same architecture for the shadow model as the target model to closely mimic its behavior, and we employ a logistic regression classifier as the final attack model.
  4. (4)
    Likelihood Ratio Attack (LiRA): We also evaluate the Likelihood Ratio Attack (LiRA) (Carlini et al., 2022), a state-of-the-art membership inference attack. Unlike threshold-based or shadow-classifier approaches, LiRA formulates membership inference as a per-sample likelihood test. Concretely, we first train multiple shadow models on disjoint subsets of data sampled from the same distribution as the target model. For each shadow model, we record the per-sample cross-entropy losses on both member (IN) examples, which are those used in training the shadow model, and non-member (OUT) examples, those excluded from training. Gaussian distributions are then fitted to these two sets of losses, yielding estimates of $P(\ell\mid\text{IN})$ and $P(\ell\mid\text{OUT})$, where $\ell$ denotes the loss.
    For each target sample, we compute the log-likelihood ratio
    $\text{LLR}(x)=\log\frac{P(\ell(x)\mid\text{IN})}{P(\ell(x)\mid\text{OUT})},$
    and classify membership based on the calibrated sign of this statistic. Importantly, post-hoc defenses such as DynaNoise are applied only when evaluating the target model, while the attacker’s shadow models are always trained without defenses. In our setup, we train five shadow models per dataset (same architecture as the target), and use their pooled losses to fit the Gaussian parameters and calibrate the decision rule.
Rationale for attack selection: The attacks above were chosen to cover the major classes of membership inference strategies studied in the literature. The confidence-threshold and loss-threshold attacks represent metric-based MIAs that exploit scalar statistics derived from model outputs, such as maximum confidence or per-sample loss. Entropy-based and modified entropy attacks further extend this class by leveraging global properties of the output distribution, capturing overconfidence patterns that are not visible through single-score metrics. Shadow-model attacks represent the class of supervised attacker training, in which the adversary trains an auxiliary classifier to explicitly learn a member versus non-member decision rule from model outputs. Finally, LiRA is a state-of-the-art adaptive statistical attack that performs per-sample likelihood testing using calibrated shadow models and is known to substantially outperform earlier methods, particularly at low false-positive rates.
Each membership inference attack is systematically evaluated under the following six defense conditions:
@@
    Adversarial Regularization (AdvReg) (Nasr et al., 2018) Defense: The model is trained with an auxiliary adversary that penalizes leakage of membership information during optimization. This regularization term encourages the model to produce output distributions that are indistinguishable between members and non-members, thereby reducing attack success rates.
  3. (3)
    MemGuard (Jia et al., 2019) defense: MemGuard generates small adversarial perturbations to output probabilities at inference time, reducing membership leakage without modifying or retraining the underlying model.
  4. (4)
    RelaxLoss (Chen et al., 2022) defense: The RelaxLoss defense is applied during model training, where a modified loss function penalizes overconfident predictions to reduce overfitting. This adjustment smooths the model’s decision boundaries and lowers membership leakage, providing an effective privacy-utility trade-off without requiring additional noise or architectural changes.
  5. (5)
    SELENA (Tang et al., 2022) defense: The SELENA defense approach is employed, which leverages ensemble learning and self-distillation to mitigate membership leakage prior to the execution of the attack.
  6. (6)
    HAMP (Chen and Pattabiraman, 2024) defense: The HAMP defense is applied by mitigating model overconfidence through entropy-regularized training with high-entropy soft labels, followed by a testing-time output modification that uniformly suppresses prediction confidence in a query-independent manner.
@@ 7.4. Experimental Setup and Results
We conduct experiments on CIFAR-10, ImageNet-10, and SST-2 datasets. For each dataset, 70% of the data is used to train and test the target model, while the remaining 30% is allocated for training shadow and attack models. All models are trained for 30 epochs using stochastic gradient descent (SGD) with a learning rate of 0.01 and a batch size of 64. Experiments are conducted on a machine equipped with NVIDIA RTX 5000 GPU to ensure efficient training and evaluation.
As our experimental baseline defenses, we implement Adversarial Regularization (AdvReg) (Nasr et al., 2018), RelaxLoss (Chen et al., 2022), MemGuard (Jia et al., 2019), SELENA (Tang et al., 2022), and HAMP (Chen and Pattabiraman, 2024). For AdvReg, we adopt the standard min-max training setup with a regularization weight $\lambda_{\text{reg}}=3.0$, one adversarial step per batch ($k_{\text{attack}}=1$), and learning rates of $10^{-2}$ and $10^{-3}$ for the classifier and attacker, respectively. For RelaxLoss, we follow the loss-smoothing formulation in the original paper and use its key hyperparameters, namely the cross-entropy threshold $\alpha=0.5$ and learning rate $lr=0.01$. For MemGuard (Jia et al., 2019) we follow the original paper’s configuration, using $\varepsilon=0.5$ and Phase-I coefficients $(c_{2}=10,\ c_{3}=0.1)$, with a defense-classifier trained for $400$ epochs. For SELENA, we use the configuration described in Section 2.2, with $K=25$ sub-models and $L=10$ partitions per sample. Finally, for HAMP, we follow Chen et al. (Chen and Pattabiraman, 2024), which enforces less-confident predictions using high-entropy soft labels and an entropy-based regularizer during training, and applies the testing-time output modification described in the original work. We use the authors’ default parameters: the entropy threshold $\gamma=0.95$ for generating high-entropy soft labels and the regularization strength $\alpha=0.001$ in the training objective.
@@
Each target model is evaluated against multiple types of membership inference attacks, including Confidence Threshold Attack, Loss Threshold Attack, Entropy-based Attack, M-Entropy Attack, Shadow Model Attack, and the Likelihood Ratio Attack (LiRA). For each attack, we report the Attack Success Rate (ASR) under seven conditions: (i) without any defense (None), (ii) AdvReg, (iii) MemGuard, (iv) RelaxLoss, (v) SELENA, (vi) HAMP, and (vii) our proposed approach DynaNoise.
@@
To quantify the privacy-utility trade-off, we utilize our proposed metric called MIDPUT, as described in Section 7.1. MIDPUT measures how effectively a defense reduces membership inference attack success rates while preserving the predictive accuracy of the target model. We report both the overall MIDPUT score and per-attack values ($\text{MIDPUT}_{Conf}$, $\text{MIDPUT}_{Loss}$, $\text{MIDPUT}_{Ent}$, $\text{MIDPUT}_{M\text{-}Ent}$, $\text{MIDPUT}_{Shadow}$, and $\text{MIDPUT}_{LiRA}$) for each evaluated defense.
Tables 3 and 4 summarize the experimental results across all datasets under a diverse set of defense mechanisms. Overall, DynaNoise demonstrates a consistently strong balance between privacy protection and model utility, achieving substantial reductions in membership inference attack success rates while preserving near-baseline target model accuracy across both vision and NLP workloads. In contrast to several prior defenses that obtain privacy gains primarily through aggressive output suppression or retraining-based regularization, DynaNoise maintains stable performance across datasets and attack types by adapting the magnitude of noise to the sensitivity of each individual query.
On CIFAR-10, DynaNoise achieves the highest overall MIDPUT score across all evaluated attacks, outperforming training-time defenses such as AdvReg and RelaxLoss, post-processing approaches such as MemGuard, and the ensemble-based SELENA method. Although HAMP reduces all attack success rates to chance level, it does so by noticeably degrading target model accuracy, resulting in near-zero net MIDPUT improvement. This outcome highlights a limitation of global confidence suppression on vision models, where uniformly flattening prediction distributions can unnecessarily distort low-risk queries and lead to avoidable utility loss.
A similar behavior is observed on ImageNet-10. HAMP again attains low attack success rates under strong attacks, including Shadow and LiRA, but at the cost of severe accuracy degradation, yielding the lowest MIDPUT scores among all evaluated defenses. Adversarial Regularization, RelaxLoss, and MemGuard provide only marginal or inconsistent privacy-utility improvements, while SELENA offers moderate protection but suffers from reduced accuracy and nontrivial computational overhead. In contrast, DynaNoise preserves near-baseline accuracy while consistently reducing attack success rates, resulting in the highest overall MIDPUT score on this dataset and demonstrating robustness under strong membership inference attacks.
On SST-2, where membership signals are weaker and modern pretrained language models exhibit relatively high-entropy and well-calibrated predictions, the behavior of defenses differs from that observed on vision datasets. In this setting, HAMP attains the highest overall MIDPUT score by reducing attack success rates to near-chance levels while slightly improving target model accuracy, which aligns well with its global confidence smoothing strategy in an already well-calibrated prediction setting. DynaNoise also substantially reduces attack success rates and preserves near-baseline accuracy, but achieves lower overall MIDPUT values than HAMP on this dataset. Notably, DynaNoise does so without uniform confidence suppression, instead selectively injecting noise only for low-entropy, higher-risk predictions.
A notable pattern is particularly evident on SST-2: Several training-based defenses such as Adversarial Regularization and MemGuard provide little to no privacy-utility improvement, yielding near-zero or negative MIDPUT gains. This behavior is consistent with the design assumptions discussed in the original AdvReg (Nasr et al., 2018) and MemGuard (Jia et al., 2019) works, which rely on pronounced separability between member and non-member predictions to supply a meaningful adversarial signal. For modern pretrained NLP models such as DistilBERT, member and non-member outputs are already highly similar and well calibrated, leaving limited room for these defenses to operate effectively. In contrast, defenses that act directly on prediction calibration, including RelaxLoss, SELENA, HAMP, and DynaNoise, remain effective in this setting.
@@
Table 3. Model accuracy and MIA success rates on different datasets under various defense mechanisms.